OpenStudy (anonymous):

Solve for x: 2x – 1 = x – 6 A. x = 4 B. x = -4 C. x = -5 D. x = 5

OpenStudy (anonymous):

(2x-x)=-6+1 x = -5 C is the correct answer

OpenStudy (anonymous):

Thanks(: